OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of 2/3

OpenStudy (campbell_st):

well knowing the slope you can set the equation up in slope intercept form as y = 2/3 x + b substitute x = 6 and y = 3 to find b then write the equation in standard form

OpenStudy (anonymous):

but how do you do that?

OpenStudy (campbell_st):

so what value did you get for b..?

OpenStudy (anonymous):

i know it would be 3=2/3(6)+b but idk how to do the rest

OpenStudy (campbell_st):

3 = 2/3 * 6 + b so 3 = 4 + b whats the value of b..?

OpenStudy (anonymous):

b=-1

OpenStudy (campbell_st):

great so the equation is y = 2/3x - 1 next multiply each term by 3 what do you get..? you are wanting to get rid of the fraction.

OpenStudy (anonymous):

3y=2x-1

OpenStudy (anonymous):

or wait no 3y=2x-3

OpenStudy (campbell_st):

a good try, but you forgot about -1 3y = 2x - 3 the standard form of a line is Ax +By = C so subtract 2x from both sides of the equation, what do you get?

OpenStudy (anonymous):

-3=3y-2x -2x+3y=-3

OpenStudy (campbell_st):

yep the both are correct and in standard form. The common rule is to have A as positive so multiply each term by -1 and you get 2x - 3y = 3 or you can have -2x +3y = -3 either is fine hope its helped

OpenStudy (anonymous):

yes thank you!